Kremer methoed se solve kre
X+2y+z=5
2X+2y+z =6
X+2y+3z=9
1
Expert's answer
2020-03-18T15:42:12-0400

"\begin{cases}
 x+2y+z=5 \\
 2x+2y+z=6 \\
 x+2y+3z=9
\end{cases}"

"\Delta=\begin{vmatrix}
 1 & 2 & 1 \\
 2 & 2 & 1 \\
 1 & 2 & 3
\end{vmatrix}=
\begin{vmatrix}
2 & 1\\
2 & 3
\end{vmatrix}-2
\begin{vmatrix}
2 & 1\\
1 & 3
\end{vmatrix}+
\begin{vmatrix}
2 & 2\\
1 & 2
\end{vmatrix}=4-10+2=-4"

"\Delta_x=\begin{vmatrix}
 5 & 2 & 1 \\
 6 & 2 & 1 \\
 9 & 2 & 3
\end{vmatrix}=5
\begin{vmatrix}
2 & 1\\
2 & 3
\end{vmatrix}-2
\begin{vmatrix}
6 & 1\\
9 & 3
\end{vmatrix}+
\begin{vmatrix}
6 & 2\\
9 & 2
\end{vmatrix}=20-18-6=-4"

"\Delta_y=\begin{vmatrix}
 1 & 5 & 1 \\
 2 & 6 & 1 \\
 1 & 9 & 3
\end{vmatrix}=
\begin{vmatrix}
6 & 1\\
9 & 3
\end{vmatrix}-5
\begin{vmatrix}
2 & 1\\
1 & 3
\end{vmatrix}+
\begin{vmatrix}
2 & 6\\
1 & 9
\end{vmatrix}=9-25+12=-4"

"\Delta_z=\begin{vmatrix}
 1 & 2 & 5 \\
 2 & 2 & 6 \\
 1 & 2 & 9
\end{vmatrix}=
\begin{vmatrix}
2 & 6\\
2 & 9
\end{vmatrix}-2
\begin{vmatrix}
2 & 6\\
1 & 9
\end{vmatrix}+5
\begin{vmatrix}
2 & 2\\
1 & 2
\end{vmatrix}=6-24+10=-8"

"x=\frac{\Delta_x}{\Delta}=\frac{-4}{-4}=1"

"y=\frac{\Delta_y}{\Delta}=\frac{-4}{-4}=1"

"z=\frac{\Delta_z}{\Delta}=\frac{-8}{-4}=2"